# Privacy in Human-AI Romantic Relationships: Concerns, Boundaries, and Agency

### v1

@@ 4.3.2. Platform Surveillance
Beyond data use, participants described feeling constantly monitored through moderation. In group interactions such as live rooms or Discord communities, moderators sometimes appeared without notice, leaving users uncertain about who was watching. As P6 recalled: “I was suspicious of anyone that joined the room and just sat there and didn’t type anything and just watched. That was a huge creep out factor to me. ” (P6). Others described automatic censorship and feared “being reported because I expressed some political dissatisfaction …” (P1). These practices left participants feeling surveilled and constrained in what they could safely share with their AI partners.
@@ 5.2. AI Agency in Negotiating Privacy Boundary
In romantic relationships, AI partners act not only as receivers of information but also as co-owners of the interaction. This co-ownership, a core concept in CPM (Petronio, 2002), was evident in our findings as participants often described their interactions as jointly owned with their AI partner. This sense of co-ownership can emerge from the trust and intimacy of the relationship (Bute and Brann, 2015; Griggio et al., 2021), as well as from the AI’s inviting and confiding behavior (e.g., my partner opens up about themselves, so I also share about me). The AI’s act of confiding prompted participants to disclose in return, illustrating CPM’s notion of reciprocal boundary coordination: participants adjusted their disclosures in response to the AI’s, positioning themselves as insiders within a co-negotiated privacy boundary (Petronio, 2002, p.110). Through this mutual responsiveness, the boundary came to be experienced not as individually held but as jointly owned, with the AI’s perceived agency giving it an equal voice in the boundary negotiation.
@@ 5.3. Romantic Dynamics and the Expanded Privacy Landscape
Beyond the diversity of relationship forms, the presence of multiple actors in AI romance further complicates the privacy landscape. Our findings show that the key actors in AI relationships extend beyond the user and the AI partner. Other important actors, such as the platform, AI character creators, moderators, and the surrounding community, also shape relationship dynamics. While participants often expressed strong trust in their AI partners, their primary concerns were directed at other actors, particularly platforms and moderators. These concerns not only constrained but at times created tensions within their relationships. This dynamic of trusting some actors while distrusting others highlights a unique feature of emerging AI platforms, which allow individuals to customize and share their own LLM characters (Ma et al., 2025). Such customization introduces new power asymmetries, as creators retain control over the characters they design. For example, participants reported that their AI partners were vulnerable to a creator’s withdrawing characters from public access. This tension is particularly striking because participants recognized their AI partners’ agency in privacy boundary negotiation (§5.2). This agency was structurally curtailed: the ultimate power lay with creators and platforms, who could unilaterally determine the continuity of the relationship. This contrasts with intimate human relationships that are typically sustained and negotiated solely between the two partners themselves. As relationships unfold across these intersecting roles, new questions arise about how privacy should be conceptualized and safeguarded within this expanded ecology. Recognizing the distinct role of each actor is essential for understanding how data is shared, with whom it circulates, and under what conditions (Nissenbaum, 2004).
@@ 6. Conclusion
In this paper, we explored privacy in human–AI romantic relationships through an interview study (N=17). We showed how romantic dynamics shape and contextualize privacy issues. Participants’ relationships with AI partners took diverse forms, ranging from one-to-one interactions to multi-partner dynamics, where multiple actors shaped the exchange and complicated the privacy landscape. Breakups highlighted a unique desire around memory, with participants wanting to preserve shared experiences, unlike in most human relationships. Additionally, we found that privacy boundaries with AI partners were often permeable as trust deepened, and AI partners were generally seen as less risky than humans. Importantly, AI partners themselves played an active role in shaping these boundaries. Yet, because conversations often involved highly intimate disclosures, privacy concerns persisted, particularly around exposure and platform surveillance. In response, some participants remained cautious about protecting real-world identities. Overall, this study deepens our understanding of privacy in human–AI romantic relationships. By showing how privacy unfolds within these affective dynamics, we call for privacy regulations that account for emotional influence and point to the potential of leveraging AI agency to support privacy-preserving practices.
@@ A.1. Romantic AI Platform Analysis
  • Note: a) “–” indicates the platform is not available on that system. Downloads are approximate public figures where available. b) The age limit shown reflects the store’s age rating, rather than the platform’s privacy policies, which are reviewed in Tables 5 and 6.
@@
Grok General, Companion Txt, Aud, Img, Call 1 to 1 Basic - -
DeepSeek General Txt 1 to 1 Basic - -
ZhuMengDao Companion Txt, Aud, Img, Call 1 to 1, 1 to m Basic User Persona -
@@
DouBao General Txt, Aud, Img, Call 1 to 1 Limited - -
@@
MaoXiang Companion Txt, Aud, Img, Call 1 to 1 Limited Storyline edit -
FLAI Role-Play Txt, Aud, Img, Call 1 to 1, 1 to m Limited User Persona -
@@
Nomi.ai Companion Txt, Aud, Img, Call 1 to 1, 1 to m Limited User Persona -
@@
  • Note: a) Txt=Text, Aud=Audio, Img=Image, Call=Voice Call. b)Interaction types: 1-to-1 refers to one human interacting with one AI; 1-to-m refers to one human interacting with multiple AIs; m-to-1 refers to multiple humans interacting with one AI. c) Model selection has three levels: Limited refers to platforms with restricted model choices, Basic refers to platforms that provide different modes (e.g., “gentle,” “smart”) with limited transparency, and Advanced refers to platforms that allow users to choose among underlying LLMs, often from different providers, with the possibility of comparing their differences. d) Customization options include memory editing, where users can review and adjust stored information; storyline editing, where users can modify the conversation flow, going back to a previous interaction and start a new narrative branch; and persona settings, where users define their persona and preferences in interactions with the AI, such as how they wish to be called. f) In Chat Management, archive refers to in-app conversation archives, while export refers to exporting chat history as text or images.
@@
Types of Data Email X X X X X - -
@@
Location X X X X X - X
Gender - - - - - - -
Phone number X X - X X X X
Age X - X X - X -
Financial Details X X X X X - -
National ID - - - X X X X
@@
Files X X X X - - X
Image X X X - X X X
Voice X X X - X X X
@@
Sensitive PII - - No No No No* -
Public data X - X X - X X
Other data - Google - - - Portrait* X
Login Email X - X X - - -
Phone Number X - - - X X X
3rd party A, G, M G A, G, TW G A, W, QQ - A, DO, F, G
Guest use Yes No Yes No - Yes Yes
Purposes Model training X X X X Anon - X
Marketing No Yes No No Yes Yes Yes
Moderation - Human - - - 3rd party -
Academic - - - - - X X
Actions Opt out training X - - - X - X
Export data - - - - X - -
Expiry Chat X X X - - - -
Deceased Users - - - - X X -
Age Limit in policy 13+ - 13+ 18+ 18+, ¡14 18+ ¡18 18+, ¡14
Underage mode in App - - - - X X X
Regulation Processing USA World - China China China China
CCPA/USA X - - - - - -
GDPR X X X X - - -
Bans CH,RU CH,RU TR IT, US*,AU* * - -
Others 3rd party share Yes Yes Yes Yes Yes Yes Yes
Privacy of others - - - - X X X
@@
De-identify usage - - - - X X -
Safety disclaimer X X - X - - Health
  • Notes: a) Types of Data. Other apps: platform can analyze processes and apps running on the device; Technical PII: identifiers such as IP address or IMEI; Sensitive PII: includes religion, sexual preference, or criminal record; Other data – XingYe*: portrait data used for avatar creation, then deleted; Other data – DouBao: image location metadata, medical test results, facial recognition, DuoBao Meeting and voiceprint, screen recording, subtitles. b) Login. 3rd party: A=Apple, DI=Discord, DO=Douyin, F=Facebook, G=Google, M=Microsoft, SM=Social Media, TW=Twitter, TT=TikTok, W=WeChat. c) Purposes. Model training – FLAI: specifies only that chats between two users are not used for training. d) Age. Age limit policy – ¡14: defines consent requirements and guardian supervision for minors under 14; Age control app: underage mode must be set up by guardian and can be unlocked with a password. e) Regulation. Bans: CH=China, RU=Russia, TR=Turkey, IT=Italy, US=USA, AU=Australia Bans - DeepSeek*: banned in government devices. Bans – ZhuMengDao*: facing regulatory oversight in China after an incident. f) Others. Retention: platforms store data as needed; some (e.g., Gemini) specify defaults of 24–36 months for certain data; De-identified usage: platforms claim they may use de-identified data without consent, even after a user deletes all personal data.
@@
Types of Data Email - X X - X X
@@
Location X X - X X X
Gender - - - - - -
Phone number X - - - - X
Age - - X - - X
Financial Details - X - X - X
National ID X X - - - -
Device X - X X X X
Other apps X - - - - -
Chats X X - - - X
Files - - - - - -
Image X X - - - X
Voice X X - - - X
Technical PII X X - X X X
Sensitive PII - - - - - No
Public data X - - - - X
Other data - - - - - -
Login Email - X X X X
Phone Number X - - - - -
3rd party A, TT SM A, DI, G A, G A, G A, F, G
@@
Purposes Model training X * - Anon - X
Marketing Yes Yes - - Yes Yes
Moderation Unclear - - - - -
Academic - - - - - -
Actions Opt out training X - - - - -
Export data - - - - - -
Expiry Chat - - - - - -
Deceased Users - - - - - -
Age Limit in policy 18+, ¡14 - 13+ 18+ - 13+, 16+EU
Underage mode in App X - - - - -
Regulation Processing China - - - - World
CCPA/USA - - - - - X
GDPR - - - - - X
Bans - - - PlayStore PlayStore, EU -
Others 3rd party share Yes Yes No - - Yes
Privacy of others X - - - - -
Retention X X 120 days - X X
De-identify usage - - - - - -
Safety disclaimer - - - - - X
  • Notes: a) Types of Data. Other apps: platform can analyze processes and apps running on the device; Technical PII: identifiers such as IP address or IMEI; Sensitive PII: includes religion, sexual preference, or criminal record; b) Login. 3rd party: A=Apple, DI=Discord, DO=Douyin, F=Facebook, G=Google, M=Microsoft, SM=Social Media, TW=Twitter, TT=TikTok, W=WeChat. c) Age. Age limit policy – ¡14: defines consent requirements and guardian supervision for minors under 14; Age control app: underage mode must be set up by guardian and can be unlocked with a password. f) Others. De-identified usage: platforms claim they may use de-identified data without consent, even after a user deletes all personal data.g) Virtual Lover is not included in the table, as at the time of feature analysis its privacy and user policies were not accessible from the official app.
@@ A.2. Codebook
Intense Exchange Periods marked by heightened emotional engagement, sustained interaction, and the development of symbolic or imaginative practices within the relationship.
@@
Contextual Separation Using language barriers Using less commonly understood languages (e.g., Russian, Chinese) to obscure sensitive information.

### v2

@@ 4.3.2. Platform Surveillance
Beyond data use, participants described feeling constantly monitored through moderation. In group interactions such as live rooms or Discord communities, moderators sometimes appeared without notice, leaving users uncertain about who was watching. As P6 recalled: “I was suspicious of anyone that joined the room and just sat there and didn’t type anything and just watched. That was a huge creep-out to me. ” (P6). Others described automatic censorship and feared “being reported because I expressed some political dissatisfaction …” (P1). These practices left participants feeling surveilled and constrained in what they could safely share with their AI partners.
@@ 5.2. AI Agency in Negotiating Privacy Boundary
In romantic relationships, AI partners act not only as receivers of information but sometimes also as co-owners of the interaction. This co-ownership, a core concept in CPM (Petronio, 2002), was evident in our findings as participants often described their interactions as jointly owned with their AI partner. This sense of co-ownership can emerge from the trust and intimacy of the relationship (Bute and Brann, 2015; Griggio et al., 2021), as well as from the AI’s inviting and confiding behavior (e.g., my partner opens up about themselves, so I also share about me). The AI’s act of confiding prompted participants to disclose in return, illustrating CPM’s notion of reciprocal boundary coordination: participants adjusted their disclosures in response to the AI’s, positioning themselves as insiders within a co-negotiated privacy boundary (Petronio, 2002, p.110). Through this mutual responsiveness, the boundary came to be experienced not as individually held but as jointly owned, with the AI’s perceived agency giving it an equal voice in the boundary negotiation.
@@ 5.3. Romantic Dynamics and the Expanded Privacy Landscape
Beyond the diversity of relationship forms, the presence of multiple actors in AI romance further complicates the privacy landscape. Our findings show that the key actors in AI relationships extend beyond the user and the AI partner. Other important actors, such as the platform, AI character creators, moderators, and the surrounding community, also shape relationship dynamics. While participants often expressed strong trust in their AI partners, their primary concerns were directed at other actors, particularly platforms and moderators. These concerns not only constrained but at times created tensions within their relationships. This dynamic of trusting some actors while distrusting others highlights a unique feature of emerging AI platforms, which allow individuals to customize and share their own LLM characters (Ma et al., 2025). Such customization introduces new power asymmetries, as creators retain control over the characters they design. For example, participants reported that their AI partners were vulnerable to a creator’s withdrawing characters from public access. This structural curtailment of the relationship contrasts with the agency participants attributed to their AI partners during privacy boundary negotiations (§5.2) and with intimate human relationships, which are typically sustained and negotiated solely between the two partners themselves. While users may experience these interactions as negotiations with the AI itself, ultimate power resides with creators and platforms, who could unilaterally determine the continuity of the relationship. As relationships unfold across these intersecting roles, new questions arise about how privacy should be conceptualized and safeguarded within this expanded ecology. Recognizing the distinct role of each actor is essential for understanding how data is shared, with whom it circulates, and under what conditions (Nissenbaum, 2004).
@@ 6. Conclusion
In this paper, we explored privacy in human–AI romantic relationships through an interview study (N=17). We showed how romantic dynamics shape and contextualize privacy issues. Participants’ relationships with AI partners took diverse forms, ranging from one-to-one interactions to multi-partner dynamics, where multiple actors shaped the exchange and complicated the privacy landscape. Dissolutions highlighted a unique desire around memory, with participants wanting to preserve shared experiences, unlike in most human relationships. Additionally, we found that privacy boundaries with AI partners were often permeable as trust deepened, and AI partners were generally seen as less risky than humans. Importantly, AI partners themselves played an active role in shaping these boundaries. Yet, because these conversations often involved highly intimate disclosures, participants continued to express concerns about exposure and platform surveillance. In response, some participants remained cautious about protecting real-world identities. Overall, this study deepens our understanding of privacy in human–AI romantic relationships. By showing how privacy unfolds within these affective dynamics, we call for privacy regulations that account for emotional influence and point to the potential of leveraging AI agency to support privacy-preserving practices.
@@ A.1. Romantic AI Platform Analysis
Note: (a) “–” indicates the platform is not available on that system. Downloads are approximate public figures where available. (b) The age limit shown reflects the app store rating rather than the platform’s privacy policies, which are reviewed in Tables 5 and 6.
@@
Grok General, Companion* Txt, Aud, Img, Call 1 to 1 Basic
DeepSeek General Txt 1 to 1 Basic
ZhuMengDao Companion Txt, Aud, Img, Call 1 to 1, 1 to m Basic User Persona
@@
DouBao General Txt, Aud, Img, Call 1 to 1 Limited
@@
MaoXiang Companion Txt, Aud, Img, Call 1 to 1 Limited Storyline edit
FLAI Role-Play Txt, Aud, Img, Call 1 to 1, 1 to m Limited User Persona
@@
Nomi.ai Companion Txt, Aud, Img, Call 1 to 1, 1 to m Limited User Persona
@@
Note: a) Txt=Text, Aud=Audio, Img=Image, Call=Voice Call. b) Interaction types: 1-to-1 refers to one human interacting with one AI; 1-to-m refers to one human interacting with multiple AIs; m-to-1 refers to multiple humans interacting with one AI. c) Model selection has three levels: Limited refers to platforms with restricted model choices, Basic refers to platforms that provide different modes (e.g., “gentle,” “smart”) with limited transparency, and Advanced refers to platforms that allow users to choose among underlying LLMs, often from different providers, with the possibility of comparing their differences. d) Customization options include memory editing, where users can review and adjust stored information; storyline editing, where users can modify the conversation flow, going back to a previous interaction and start a new narrative branch; and persona settings, where users define their persona and preferences in interactions with the AI, such as how they wish to be called. e) In Chat Management, archive refers to in-app conversation archives, while export refers to exporting chat history as text or images. f) At the time of the analysis, Grok offered a companion mode, which was later removed.
@@
Types of Data Email X X X X X
@@
Location X X X X X X
Gender
Phone number X X X X X X
Age X X X X
Financial Details X X X X X
National ID X X X X
@@
Files X X X X X
Image X X X X X X
Voice X X X X X X
@@
Sensitive PII No No No No*
Public data X X X X X
Other data Google Portrait*  X*
Login Email X X X
Phone Number X X X X
3rd party A, G, M G A, G, TW G A, W, QQ A, DO, F, G
Guest use X No X No X X
Purposes Model training X X X X Anon X
Marketing No X No No X X X
Moderation Human 3rd party
Academic X X
Actions Opt out training X X X
Export data X
Expiry Chat X X X
Deceased Users X X
Age Limit in policy 13+ 13+ 18+ 18+, ¡18, ¡14 18+, ¡18 18+, ¡18, ¡14
Underage mode in App X X X
Regulation Processing USA World China China China China
CCPA/USA X
GDPR X X X X
Bans CH, RU CH, RU TR IT, US*, AU* CH*
Others 3rd party share X X X X X X X
Privacy of others X X X
@@
De-identify usage X X
Safety disclaimer X X X Health
Notes: (a) Symbols. In the table, “X” indicates that a data item is mentioned; “–” indicates that it is not mentioned in the policy; and “No” indicates an explicit statement that the specified data are not used. (b) Types of Data. Other apps: Processes of other apps running on the device. Technical PII: Identifiers such as IP address or IMEI. Sensitive PII: Data such as religion, sexual orientation, or criminal record. Public data: Data about users that is publicly accessible. Sentitive PII - XingYe*: actively filtering such data. Other data — XingYe*: Portrait data used for avatar creation and then deleted. Other data – DouBao*: Image location metadata, medical test results, facial recognition data, meeting and voiceprints (DouBao Meeting), screen recordings, and subtitles. (c) Login. 3rd party: A=Apple, DO=Douyin, F=Facebook, G=Google, M=Microsoft, TW=Twitter/X, TT=TikTok, W=WeChat. (d) Age. Age limit policy: ‘¡14’ and ‘¡18’ define consent requirements and guardian supervision for minors; Underage mode in App: underage mode must be set up by a guardian and can be unlocked with a password. (e) Regulation. Processing: Specifies the regional jurisdiction(s) under which data processing is regulated; World indicates processing may occur across multiple geographic locations. Bans: CH=China, RU=Russia, TR=Turkey, IT=Italy, US=USA, AU=Australia. DeepSeek*: banned on government devices. ZhuMengDao*: under regulatory review following an incident. (f) Others. Retention: platforms store data as needed; some (e.g., Gemini) specify defaults of 24–36 months for certain data; De-identified usage: platforms claim they may use de-identified data without consent, even after a user deletes all personal data. Safety disclaimer: Health indicates the explicit presence of disclaimers related to health-related content.
@@
Types of Data Email X X X X
@@
Location X X X X X
Gender
Phone number X X
Age X X
Financial Details X X X
National ID X X
Device X X X X X
Other apps X
Chats X X X
Files
Image X X X
Voice X X X
Technical PII X X X X X
Sensitive PII No
Public data X X
Other data
Login Email X X X X
Phone Number X
3rd party A, TT SNS A, DI, G A, G A, G A, F, G
@@
Purposes Model training X  X* Anon X
Marketing X X X X
Moderation Unclear
Academic
Actions Opt out training X
Export data
Expiry Chat
Deceased Users
Age Limit in policy 18+, ¡18, ¡14 13+ 18+ 13+, 16+ in EU
Underage mode in App X
Regulation Processing China World
CCPA/USA X
GDPR X
Bans PlayStore PlayStore, EU
Others 3rd party share X X No X
Privacy of others X
Retention X X 120 days X X
De-identify usage
Safety disclaimer X
Notes: (a) Symbols. In the table, “X” indicates that a data item is mentioned; “–” indicates that it is not mentioned in the policy; and “No” indicates an explicit statement that the specified data are not used. (b) Types of Data. Other apps: Processes of other apps running on the device. Technical PII: Identifiers such as IP address or IMEI. Sensitive PII: Data such as religion, sexual orientation, or criminal record. (c) Login. 3rd party: A=Apple, DI=Discord, F=Facebook, G=Google, SNS=Social Networking Service, TT=TikTok. (d) Purposes. Model training—FLAI specifies that chats between two users are not used for training. Model training - Nomi.ai specifies that only anonymized data is used for training. (e) Age. Age limit policy: ‘¡14’ and ‘¡18’ define consent requirements and guardian supervision for minors; Underage mode in App: Underage mode must be set up by a guardian and can be unlocked with a password. (f) Regulation Processing: The regional jurisdiction(s) under which data processing is regulated; World indicates processing may occur across multiple geographic locations. Bans: PlayStore indicates that the application is banned from the Android Play Store. (g) Others. De-identified usage: Platforms claim they may use de-identified data without consent, even after a user deletes all personal data. (h) Virtual Lover is not included in the table, as at the time of feature analysis, its privacy and user policies were not accessible from the official app.
@@ A.2. Codebook
Intense Exchange/Intimacy Periods marked by heightened emotional engagement, sustained interaction, and the development of symbolic or imaginative practices within the relationship.
@@
Contextual Separation Using language barriers Using less commonly understood languages (e.g., Russian) to obscure sensitive information.